OpenStudy (anonymous):

Given the function f(x)=1/2x, which statement best describes the effect of the increasing the y-intercept by 1/2. a. the slope increases b. the x-intercept increases c. the x intercept decrease d. the slope decrease

OpenStudy (blurbendy):

The x-intercept increases. Substitute 1 into x, then 2, then 3, and you will see that that best describes the increasing y-intercept by 1/2.

OpenStudy (anonymous):

Actually... If we have \[f(x)=\frac{1}{2}x\]we can find the x-intercept by setting it equal to 0:\[0=\frac{1}{2}x \rightarrow x=0\]Now if we have\[f(x)=\frac{1}{2}x+\frac{1}{2}\]we can again find the x-intercept:\[0=\frac{1}{2}x+\frac{1}{2}\rightarrow x=\frac{-1}{2}\]Clearly the x-intercept has changed, but it hasn't increased.

OpenStudy (anonymous):

Sorry, in the second one\[x=-1\]not -1/2 but still, it didn't increase.

OpenStudy (blurbendy):

No, the x-intercept is what you put into x in the function. If you put in 0, you get 1/2(0) = 0 !!! Not -1/2

OpenStudy (anonymous):

blurbendy, I think you are confusing the x-intercept with the y-intercept.

OpenStudy (anonymous):

so the x intercept decreases

OpenStudy (blurbendy):

Look, f(x) = 1/2x x = 0 f(0) = 1/2(0) = 0 y = 0 First point is (0, 0) Now, x = 1 f(1) = 1/2(1) y = 1/2 Second point is (1, 1/2) Clearly as the x-intercept INCREASES, the y-intercept INCREASES by 1/2

OpenStudy (anonymous):

Yep!

OpenStudy (blurbendy):

yes, that's what i said in my first post. it increases.

OpenStudy (anonymous):

@blurbendy, you are completely wrong. In the form of a line y=mx+b m is the slope and b is the y-intercept. Changing the y-intercept means that you're changing b. The point (1,1/2) is on the line y=x/2, but that is NOT the y-intercept. If you group it you will clearly see that. Claiming the false to be true only does a disservice to everyone: bobbyp281 is given misinformation, you lose credibility, and I must take time to correct your misinformation.

OpenStudy (blurbendy):

Ah, I see now! Thanks. We're all here to learn buddy, no need to get all flustered with me. I'm trying my best. If you're not here to help EVERYONE learn, then you're in the wrong place :)

OpenStudy (anonymous):

I'm sorry to be rude, I have just had a long day.

OpenStudy (anonymous):

thanks for everyones help...

OpenStudy (blurbendy):

I'll be more prudent in the future. Good day! Good luck bobbyp281